OpenStudy (anonymous):

how do i work out this divison sum 54 divided by 6464

OpenStudy (saifoo.khan):

\[\huge \frac{54}{6464}\]

OpenStudy (anonymous):

I'll just add. If the number isn't divisible by the denominator and you could still simplify it, better yet simplify it for smaller terms :)

OpenStudy (saifoo.khan):

try cancelling both sides.

OpenStudy (saifoo.khan):

Start from 2.. then 3..

OpenStudy (saifoo.khan):

\[\huge \frac{54}{6464} = \frac{27}{3232}\]

OpenStudy (anonymous):

you can simplify by dividing by 2 giving 27/3232 and theres long division if you haven't got or not allowed a calculator

OpenStudy (anonymous):

0.0083539603960396039603960396039604

OpenStudy (anonymous):

54/6464 = 27/3232 (it is here an irreducible fraction and further it is a repeating decimal i.e.) 0.008353960 where 3960 is the recurring.